A plane traveled 1012 miles to go to jacksonville and back. The trip there was with wind. It took 11 hours.The trip back took 22 hours. What is the speed of the plane in still air? What is the speed of the wind?
-----------------------------------------------------------------------------------------
I tried drawing a table, but it has been a wile since I have solved any problems. The answer I got was a repeating decimal, so I doubt very much its likelihood of being correct.

Answer by mananth(16949)

plane speed=x mph
wind speed=y mph
with wind Speed = (x+y)
against wind speed= (x-y)
t=d/r
d= 1012 downstream
1012 / (x+y)= Speed = 11
divide by 11

92 / (x+y)= 1
x+ y 92 ............1
d= against wind 1012
1012 /(x-y)= 22
divide by 22
46 /(x-y) = 1
x - y = 46.00 .............2
add up (1) & (2)
2 x = 138
/ 2
x= 69 mph speed of plane in still air
plug value of x in (1)
we get y= 23 mph speed of wind